# TransMLA: Migrating GQA Models to MLA with Full DeepSeek Compatibility and Speedup

### v4

@@ Abstract
Modern large-language models often face communication bottlenecks on current hardware rather than computational limitations. Multi-head latent attention (MLA) addresses this by compressing the key-value cache using low-rank matrices, while the Absorb operation prevents the KV cache from reverting to its original size, significantly boosting both training and inference speed. Despite the success of DeepSeek V2/V3/R1, most model vendors have heavily invested in optimizing GQA-based models and therefore lack strong incentives to retrain MLA-based models from scratch. In this paper, we introduce TransMLA, a framework that seamlessly converts any GQA-based pre-trained model (e.g., LLaMA, Qwen, Mixtral) into an MLA-based model. For the first time, our method enables direct conversion of these models into a format compatible with DeepSeek’s codebase, allowing them to fully leverage DeepSeek-specific optimizations such as vLLM and SGlang. By compressing 93% of the KV cache in LLaMA-2-7B, we achieve a 10.6x speedup with an 8K context length while maintaining meaningful output. Moreover, the model requires only 6B tokens for fine-tuning to recover comparable performance across multiple benchmarks. TransMLA provides a practical path for migrating GQA-based models to the MLA structure, and when combined with DeepSeek’s advanced optimizations—such as FP8 quantization and Multi-Token Prediction—further inference acceleration can be achieved.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have become indispensable tools for productivity. Notable models include GPT-4o OpenAI (2024), DeepSeek R1 Guo et al. (2025), Claude 3.7 Sonnet Anthropic (2024), and Gemini-2.5 Team et al. (2024a), etc. Open-source models like LLaMA-4 AI@Meta (2024), Mistral-3 Mistral (2024), Qwen-3 Qwen (2024), DeepSeek V3/R1 Liu et al. (2024a), Gemma-3 Team et al. (2024b), and Phi-4 Abdin et al. (2024) have narrowed the performance gap to closed-source alternatives. The effectiveness of LLMs primarily stems from Next Token Prediction Radford (2018); Brown et al. (2020), where tokens are predicted sequentially, with attention computed between each token and the tokens preceding it. To optimize computations, key-value pairs are cached (KV cache). However, as model sizes grow, the overhead of caching increases, leading to memory and communication bottlenecks.
@@
To address these challenges, Group-Query Attention (GQA)Ainslie et al. (2023) was introduced to group query heads, where all query heads in each group shares a single key and value head. When there is one group, it becomes Multi-Query Attention (MQA) Shazeer (2019), and when the number of groups equals the number of heads, it becomes Multi-Head Attention (MHA) Vaswani et al. (2017). Both GQA and MQA reduce KV cache requirements compared to MHA, but they sacrifice model performance at the same time. Additionally, techniques such as post-training KV cache compressionXiao et al. (2024); Liu et al. (2024b); Hooper et al. (2024); Zhang et al. (2023) result in non-standard implementations that require specialized optimizations, limiting their widespread adoption.
A pre-trained KV cache compression method, Multi-Head Latent Attention (MLA), introduced in DeepSeek V2 DeepSeek-AI (2024) and extended in DeepSeek V3 DeepSeek-AI (2024) and DeepSeek R1 Guo et al. (2025), strikes a better balance between computational efficiency and model performance. These models achieve top-tier performance while reducing the training and inference costs of LLMs to a new level. Furthermore, the DeepSeek team actively contributes to the open-source ecosystem, enabling highly optimized implementation and deployment solutions for efficient use of their models.
@@
In this paper, we first prove that MLA consistently offers higher expressive power than GQA under the same KV cache overhead, which theoretically explains the advantage of MLA. However, a practical obstacle preventing model vendors from switching to MLA is the substantial prior investment on GQA-based models. This motivates us to ask, can we seamlessly convert a GQA-based pretrained model, such as LLaMA-2&3 AI@Meta (2024) and Qwen-2.5&3 Qwen (2024), to MLA so that we can inherit the model weights and pretraining effort, rather than training MLA from scratch?
A key challenge in converting a GQA-based model to MLA is the presence of Rotary Position Embedding (RoPE) (Su et al., 2024) in each query-key head, which prevents the use of the Absorb Chang et al. (2024) operation, which is crucial in DeepSeek for switching between computation-efficient and memory-efficient modes. Inspired by DeepSeek’s Decoupled RoPE strategy, we propose a solution that concentrates the positional information in keys into a small subset of embedding dimensions. The remaining dimensions, which have minimal positional information, have their RoPE removed and are merged with the values for low-rank decompositions. The key up-projection, without RoPE, can then be absorbed into the query as in DeepSeek models.
@@
Finally, we observe that the magnitude of the keys, after removing RoPE, significantly exceeds that of the values. Directly applying PCA on both the keys and values tends to preserve the principal directions of the keys only, resulting in a loss of information from the Value and significant degradation in compressed model performance. To address this issue, we propose a Balanced Key-Value method, which first balances the norms of the keys and values and then jointly performs PCA, resulting in a notable improvement in compression performance.
@@ 2 Related Work
In large language models (LLMs), the key-value (KV) cache is essential for efficient autoregressive decoding, as it stores past key and value vectors to avoid redundant computations. However, the memory consumption of the KV cache grows linearly with the sequence length, since each new token requires storing its KV pair. This increasing memory footprint can become a bottleneck when processing very long sequences. To address this challenge, various techniques have been proposed to reduce the KV cache’s memory usage while maintaining model performance.
Group-Query Attention (GQA) Ainslie et al. (2023) improves KV cache efficiency by grouping query heads so that all queries in a group share a single key and value head. When the number of groups is one, GQA degenerates to Multi-Query Attention (MQA)Shazeer (2019). Both GQA and MQA reduce KV cache memory requirements compared to MHA but often sacrifice some model performance. Multi-Head Latent Attention (MLA), introduced in DeepSeek V2 DeepSeek-AI (2024) and extended in later versions DeepSeek-AI (2024); Guo et al. (2025), is a pretrained KV cache compression method that achieves a better trade-off between computational efficiency and model accuracy. MLA significantly reduces training and inference costs while maintaining competitive performance. Tensor Product Attention (TPA) Zhang et al. (2025) takes a different approach by dynamically factorizing activations instead of static weights, enabling an order-of-magnitude reduction in inference-time KV cache size compared to MHA models. However, new attention mechanisms like TPA generally require training models from scratch, which can be costly. TransMLA does not introduce a new attention design but acts as a bridge to transform the existing GQA-based model into a MLA-based one. This transformation only requires minimal fine-tuning to restore model performance.
Another approach is to optimize the KV cache of existing pre-trained models. For example, dynamic token pruning is employed by LazyLLM Fu et al. (2024), A2SF Jo and Shin (2024), and SnapKV Li et al. (2024). These methods selectively prune less important tokens from the KV cache. Sharing KV representations across layers, as in YONO Sun et al. (2024), MiniCache Liu et al. (2024c), and MLKV Zuhri et al. (2024), reduces memory by reusing the same KV cache across multiple layers. This can drastically lower memory usage and speed up inference. However, these optimization methods often require specialized configurations, making deployment difficult and limiting their widespread use. In contrast, TransMLA leverages the mature ecosystem of DeepSeek, allowing direct loading of the transformed model, making it easy to accelerate deployment across various DeepSeek-supported platforms.
@@ 4 TransMLA
In this section, we formally introduce TransMLA, beginning with two key observations:
1. Given the same amount of KV caches, MLA consistently exhibits greater expressive capacity than GQA. The full proof is provided in Appendix A. In summary, it demonstrates that any GQA model can be reformulated as an MLA model by introducing an additional projection matrix. When RoPE is present, the MLA should be expressed in the ‘absorbed’ form. However, the reverse transformation is not always possible, which implies that MLA has greater expressive power than GQA.
2. Acceleration can only be achieved when MLA operates with fewer KV caches. Although it is feasible to construct an MLA-equivalent representation of a GQA model, accelerating inference requires reducing KV caches. To address this, we transform a GQA-based model into a DeepSeek-style architecture, ensuring full compatibility with the DeepSeek inference ecosystem.
Figure 1: The pipeline for transferring GQA-based models to MLA involves applying RoRoPE to concentrate positional information in the first head, enhanced with FreqFold. KV norms are balanced before performing low-rank decomposition.
@@ 4.1 Merging All Key Heads as One
Before compressing the KV cache, it is necessary to merge all groups of key-value heads from the GQA into a single latent head. For each query head $i$, we introduce $W^{UK}_{i}\in\mathbb{R}^{d\times gd}$ with the group index $j=\left\lceil\frac{i}{h/g}\right\rceil-1$, and initialize the matrix ${W^{UK}_{i}}{[:,jd:(j+1)d]}$ to be $I_{d}$ (identity matrix of shape $d\times d$), with all other elements set to 0. (We adopt the matrix indexing notation used in Python and PyTorch, and will continue to do so throughout this work without further explanation.) This initialization allows the projection of the key’s latent representation to be simultaneously mapped onto multiple query heads, with only the corresponding key head being multiplied by the appropriate query head. Similarly, during the computation of the multiple heads of the values, the attention scores for each head are multiplied accordingly. To keep the output unchanged, we similarly initialize $W^{UV}_{i}$ so that only the corresponding value head is an identity mapping, while all other elements are set to zero. Since we have now merged all the key heads into a single head, in order to ensure that this transformation is equivalent to the original, we also merge the RoPE operations from different heads into one large RoPE operation, denoted as $\widehat{\text{RoPE}}$, applied to the entire merged key head. Since the original RoPE operation in GQA is the same for each head, $\widehat{\text{RoPE}}$ simply applies the same RoPE operation repeatedly for every $d$ dimensions. In this way, the computation of GQA attention is transformed into the following form:
@@ 4.2 Rotating Queries and Keys to Minimize Transformation Loss Towards Decoupled RoPE
In Equation 14, the term $\left(\mathbf{\hat{q}}^{R}_{t,i}\right)^{\top}\mathbf{\hat{k}}_{j}^{R}$ can be expressed as the sum of inner products over paired dimensions across multiple attention heads, incorporating positional information:
@@
This demonstrates that the rotational transformation $\mathbf{U}_{l}$ preserves the RoPE-based inner product structure. A detailed proof of this property is provided in the Appendix B.
@@ Appendix B Proof of RoPE Inner Product Invariance under Orthogonal Transformation
In practice, we take advantage of this property to find such orthogonal matrices $U_{l}$ that concentrate the principal components of keys into the first head. In this way, when we remove the principal components of most heads in queries and keys, we can retain most of the positional information in the first head, significantly reducing the loss caused by removing RoPE.
Specifically, we proceed as follows: First, we run the model on a calibration dataset (Wikitext-2) to obtain the keys at each layer. Then, we perform PCA on the key activations across the dimensions corresponding to different heads. The resulting matrix of eigenvectors (ordered by the magnitude of their corresponding eigenvalues, from largest to smallest) is used as the orthogonal matrix $U_{l}$ in this context. As a result, after this rotation, the principal components of the keys in each dimension are concentrated in the first few heads.

### v5

# TransMLA: MLA Is All You Need

@@ Abstract
Modern large-language models often face communication bottlenecks on current hardware rather than computational limitations. Multi-head latent attention (MLA) addresses this by compressing the key-value cache using low-rank matrices, while the Absorb operation prevents the KV cache from reverting to its original size, significantly boosting both training and inference speed. Despite the success of DeepSeek V2/V3/R1, most model providers have heavily invested in optimizing GQA-based models and, therefore, lack strong incentives to retrain MLA-based models from scratch. This paper demonstrates that MLA provides superior expressive power compared to GQA with the same KV cache overhead, thereby offering a rationale for transitioning from GQA to MLA. In addition, we introduce TransMLA, a framework that seamlessly converts any GQA-based pre-trained model (e.g., LLaMA, Qwen, Gemma, Mistral/Mixtral) into an MLA-based model. For the first time, our method enables direct conversion of these models into a format compatible with DeepSeek’s codebase, allowing them to fully leverage DeepSeek-specific optimizations such as vLLM and SGlang. By compressing 93% of the KV cache in LLaMA-2-7B, we achieve a 10.6x speedup with an 8K context length while maintaining meaningful output. Moreover, the model requires only 6B tokens for fine-tuning to recover comparable performance across multiple benchmarks. TransMLA provides a practical path for migrating GQA-based models to the MLA structure, and when combined with DeepSeek’s advanced optimizations—such as FP8 quantization and Multi-Token Prediction—further inference acceleration can be achieved.
(a) Given the same KV cache size, the expressiveness increases in the order of GQA, MLA, and MQA.
(b) TransMLA concentrates positional information into $K_{rope}$ and compresses $K_{nope}$ and $V$.
Figure 1: GQA, MLA, and MQA can be equivalently transformed in one direction, illustrating a gradual increase in expressive power. RoRoPE aggregates positional information in the first head, eliminating the need for RoPE in others. FreqFold further enhances this effect. Finally, after balancing the magnitudes of $K_{rope}$ and $V$, a joint low-rank approximation is applied to compress the KV cache.
@@ 1 Introduction
Advanced Large language models (LLMs)—GPT-4o OpenAI (2024), Claude 3.7 Sonnet Anthropic (2024), Gemini-2.5 Team et al. (2024a), LLaMA-4 AI@Meta (2024), Mistral-3 Mistral (2024), Qwen-3 Qwen (2024), DeepSeek V3/R1 Liu et al. (2024a); Guo et al. (2025), Gemma-3 Team et al. (2024b), and Phi-4 Abdin et al. (2024)—rapidly evolving frontier for both research and applications. LLMs rely on next-token prediction Radford (2018); Brown et al. (2020): tokens are generated sequentially, and self-attention is computed over all preceding tokens. To avoid redundant computation, implementations store the intermediate key–value (KV) pairs in a cache. Yet, as model and context sizes grow, the KV cache itself becomes a major bottleneck for inference.
@@
To mitigate these challenges, Group-Query Attention (GQA) Ainslie et al. (2023) groups the query heads so that every head within a group shares a single key and value head. When the number of groups is one, GQA degenerates to Multi-Query Attention (MQA) Shazeer (2019). When the number of groups equals the number of heads, it reduces to the standard Multi-Head Attention (MHA) Vaswani et al. (2017). While both GQA and MQA cut the size of the KV cache relative to MHA, they do so at the cost of model quality. Post-training KV-cache compression techniques—such as Duo-Attention Xiao et al. (2024), KiVi Liu et al. (2024b), KV-Quant Hooper et al. (2024), and H2O Zhang et al. (2023)—further shrink memory usage, but their non-standard implementations demand specialized optimizations, hindering widespread adoption.
Multi-Head Latent Attention (MLA)—introduced with DeepSeek V2 DeepSeek-AI (2024) and further refined in DeepSeek V3 DeepSeek-AI (2024) and DeepSeek R1 Guo et al. (2025)—offers a pre-trained KV-cache compression strategy that strikes an excellent balance between computational efficiency and model quality. Models equipped with MLA deliver state-of-the-art results while driving training and inference costs to new lows. Moreover, the DeepSeek team’s ongoing commitment to open-source releases provides highly optimized implementations and deployment recipes, making these advances readily accessible to the community.
@@
In this paper, we first prove that MLA consistently offers higher expressive power than GQA under the same KV cache overhead, which theoretically explains the advantage of MLA. However, a practical obstacle preventing model vendors from switching to MLA is the substantial prior investment on GQA-based models. This motivates us to ask, can we seamlessly convert a GQA-based pretrained model, such as LLaMA AI@Meta (2024) and Qwen Qwen (2024), to MLA so that we can inherit the model weights and pretraining effort, rather than training MLA from scratch?
A key obstacle to converting a GQA-based model to MLA is that every query–key head carries its own Rotary Positional Embedding (RoPE) (Su et al., 2024), blocking the Absorb operation (Chang et al., 2024) that DeepSeek uses to switch between compute- and memory-efficient modes. Borrowing from DeepSeek’s Decoupled RoPE scheme, we concentrate the positional signal in $K$ into a small subset of dimensions, $K_{\text{rope}}$. The remaining dimensions, $K_{\text{nope}}$, contain little positional content; we drop their RoPE and merge them with $V$ for low-rank decomposition. Once RoPE is isolated, the key up-projection—now RoPE-free—can be absorbed into the query projection exactly as in DeepSeek, enabling seamless MLA conversion.
@@
Finally, we found that the $\ell_{2}$-norm of $K_{\text{nope}}$ is far larger than that of $V$. If we run PCA on the concatenated matrix $\bigl{[}K_{\text{nope}};V\bigr{]}$ without adjustment, the principal components are dominated by $K_{\text{nope}}$, leading to severe information loss from the value subspace and a sharp drop in accuracy. We therefore introduce a Balanced Key–Value (BKV) procedure: we first rescale $K_{\text{nope}}$ and $V$ so that their norms match, and only then perform joint PCA. This simple normalization restores balance between the two subspaces and delivers a marked improvement in compression quality.
@@ 2 Related Work
In large language models, autoregressive decoding reuses past activations by storing their key–value (KV) pairs in a cache. Because the size of this cache grows linearly with sequence length, its memory footprint quickly becomes the limiting factor for very long contexts. Consequently, shrinking the KV cache without compromising accuracy has become a pivotal research focus, motivating a spectrum of architectural innovations and compression strategies.
Multi-Query Attention (MQA) Shazeer (2019) and Group-Query Attention (GQA) Ainslie et al. (2023) shrink the KV cache by letting every query in a group share a single key and value head. Although both schemes save memory relative to Multi-Head Attention (MHA) Vaswani et al. (2017), they usually give up some accuracy. Multi-Head Latent Attention (MLA)—introduced with DeepSeek V2 DeepSeek-AI (2024) and refined in later releases DeepSeek V3/R1 DeepSeek-AI (2024); Guo et al. (2025)—offers a more favorable trade-off, delivering near-state-of-the-art quality while cutting training and inference costs. Grouped Latent Attention (GLA) Zadouri et al. (2025) provides a parallel-friendly implementation of latent attention that further accelerates MLA inference. By contrast, Tensor Product Attention (TPA) Zhang et al. (2025) tackles the memory bottleneck by dynamically factorizing activations, slashing the runtime KV cache by an order of magnitude, but it necessitates training the model from scratch. TransMLA fills this gap: rather than proposing yet another attention variant, it converts an existing GQA model into an MLA model with only light fine-tuning, restoring accuracy while inheriting MLA’s memory and speed advantages.
Another approach is to optimize the KV cache of existing pre-trained models. For example, dynamic token pruning is employed by LazyLLM Fu et al. (2024), A2SF Jo and Shin (2024), and SnapKV Li et al. (2024). These methods selectively prune less important tokens from the KV cache. Sharing KV representations across layers, as in YONO Sun et al. (2024), MiniCache Liu et al. (2024c), and MLKV Zuhri et al. (2024), reduces memory by reusing the same KV cache across multiple layers. This can drastically lower memory usage and speed up inference. Although effective, both families of methods usually require custom kernels or runtime tweaks, complicating deployment and limiting adoption. TransMLA, by contrast, plugs directly into the mature DeepSeek ecosystem—converted checkpoints load out-of-the-box, delivering MLA-level speed-ups across every DeepSeek-supported platform.
@@ 4 TransMLA
In this section we formally present TransMLA, motivated by two observations:
1. For a fixed KV-cache budget, MLA is strictly more expressive than GQA. As proven in Appendix A (and illustrated in Figure 1(a)), any GQA layer can be rewritten as an MLA layer by introducing a single additional projection matrix. The reverse transformation is not always possible, implying that MLA subsumes GQA. When Rotary Positional Embeddings (RoPE) are present, the MLA equivalent must be expressed in the absorbed form.
2. Inference acceleration occurs only when MLA uses a smaller KV cache. Although one can build an MLA-equivalent representation of a GQA model, speedups arise only if the number of stored KV vectors is actually reduced. TransMLA therefore converts a GQA-based network into a DeepSeek-like MLA architecture, allowing the transformed model to run directly on DeepSeek’s optimized inference stack and realize the full memory–latency benefits.
Figure 2: Pipeline of RoRoPE for decoupling RoPE. Blue lines denote real-part dimensions, orange lines denote imaginary-part dimensions. When the keys from multiple heads are concatenated, permuting dimensions does not change the computation, so we gather the same dimension (i.e., the same rotational frequency) across all heads and apply joint principal-component analysis—using the identical procedure for the real and imaginary parts. For each frequency we keep a single principal component, which captures the dominant positional variation and can be represented by a standard RoPE in one attention head.
@@ 4.1 Merging All Key Heads as One
Because MLA ties all KV heads to a single latent dimension, the first step in converting a GQA layer to MLA is to merge every GQA key–value group into one latent head before any KV-cache compression is applied. For each query head $i$, we introduce $W^{UK}_{i}\in\mathbb{R}^{d\times gd}$ with the group index $j=\left\lceil\frac{i}{h/g}\right\rceil-1$, and initialize the matrix ${W^{UK}_{i}}{[:,jd:(j+1)d]}$ to be $I_{d}$ (identity matrix of shape $d\times d$), with all other elements set to 0. (We adopt the matrix indexing notation used in Python and PyTorch, and will continue to do so throughout this work without further explanation.) This initialization allows the projection of the key’s latent representation to be simultaneously mapped onto multiple query heads, with only the corresponding key head being multiplied by the appropriate query head. Similarly, during the computation of the multiple heads of the values, the attention scores for each head are multiplied accordingly. To keep the output unchanged, we similarly initialize $W^{UV}_{i}$ so that only the corresponding value head is an identity mapping, while all other elements are set to zero. Since we have now merged all the key heads into a single head, in order to ensure that this transformation is equivalent to the original, we also merge the RoPE operations from different heads into one large RoPE operation, denoted as $\widehat{\text{RoPE}}$, applied to the entire merged key head. Since the original RoPE operation in GQA is the same for each head, $\widehat{\text{RoPE}}$ simply applies the same RoPE operation repeatedly for every $d$ dimensions. In this way, the computation of GQA attention is transformed into the following form:
@@ 4.2 Rotating Queries and Keys to Minimize Transformation Loss Towards Decoupled RoPE
As illustrated in Figure 2, applying RoRoPE to the merged head removes the bulk of the positional signal from K. In Equation 14, the term $\left(\mathbf{\hat{q}}^{R}_{t,i}\right)^{\top}\mathbf{\hat{k}}_{j}^{R}$ can be expressed as the sum of inner products over paired dimensions across multiple attention heads, incorporating positional information:
@@
This demonstrates that the rotational transformation $\mathbf{U}_{l}$ preserves the RoPE-based inner product structure. Simply put, because the same rotation values (i.e., $\cos t\theta_{l}$ and $\sin t\theta_{l}$) are applied identically to each dimension of the $l$-th 2D subspace across all attention heads, any orthogonal transformation $\mathbf{U}_{l}^{\top}\mathbf{U}_{l}=\mathbf{I}$ applied to these dimensions within the same subspace leaves the inner product $\mathbf{\hat{q}}^{R^{\top}}_{t,i}\mathbf{\hat{k}}_{j}^{R}$ unchanged. However, the preceding equation reveals a critical constraint: for the inner product’s value to remain unchanged after transformation, the same orthogonal matrix $\mathbf{U}_{l}$ must be applied to both the real ($12l-$ 1) and imaginary ($2l$) components of the key vectors within each 2D subspace. For a detailed proof and our proposed solution, please refer to Appendix B.
@@ A.1 Introduction
Figure 6: Comparison of Multi-Query Attention (MQA), Group Query Attention (GQA), and Multi-Head Latent Attention (MLA). In this work, we illustrate that given the same KV cache size, the expressiveness increases in the order of GQA, MLA, and MQA. In the figure, $h,d,g$ denote the number of heads, hidden dimension of each head, and the number of groups (K/V heads) in GQA, respectively. In MQA, the head dimension is set to $gd$ to align the KV cache size with GQA and MLA. As a result, the KV cache size per token per layer for all three approaches is $2gd$
@@ Appendix B Proof of RoPE Inner Product Invariance under Orthogonal Transformation
In practice, we leverage this rotational invariance property to find a set of optimal orthogonal matrices $\{\mathbf{U}_{l}\}$ that concentrate the principal components of the key vectors into the first few attention heads. The preceding proof reveals a critical constraint: for the inner product’s value to remain unchanged after transformation, the same orthogonal matrix $\mathbf{U}_{l}$ must be applied to both the real ($12l-$ 1) and imaginary ($2l$) components of the key vectors within each 2D subspace. This requirement precludes performing separate PCA on the real and imaginary parts. We must therefore find a single rotation that is jointly optimal for both.
Specifically, we formulate this as a joint optimization problem. First, we process a calibration dataset (e.g., Wikitext-2) to collect the key activations at each layer. For each RoPE subspace $l\in\{1,\dots,d/2\}$, we obtain two collections of $n\times h$-dimensional matrices (where $n$ denotes the number of samples): the "real" parts $\{\mathbf{K}_{x,l}\}_{l}$ and the "imaginary" parts $\{\mathbf{K}_{y,l}\}_{l}$. To find a single transformation $U_{l}$ that simultaneously compresses the information from both sets into the first few heads, we proceed as follows.
Let $\sigma_{x,l}=\mathbf{K}_{x,l}^{\top}\mathbf{K}_{x,l}$ and $\sigma_{y,l}=\mathbf{K}_{y,l}^{\top}\mathbf{K}_{y,l}$ be the $h\times h$ covariance matrices of the real and imaginary key components, respectively. Our objective is to find an orthogonal matrix $\mathbf{U}_{l}$ that maximizes the variance—or energy—concentrated in the first $m$ heads after rotation. This corresponds to maximizing the trace of the top-left $m\times m$ submatrix of the summed covariance of the rotated vectors. The problem is formally stated as:
$\underset{\mathbf{U}_{l}}{\max}\quad\mathrm{Tr}\left[(\mathbf{U}_{l}^{T}(\sigma_{x,l}+\sigma_{y,l})\mathbf{U}_{l})_{:m,:m}\right]\quad\text{s.t.}\quad\mathbf{U}_{l}^{T}\mathbf{U}_{l}=I.$ (32)
Here, $\mathbf{U}_{l}$ is the $h\times h$ orthogonal optimization variable, and $(\cdot)_{:m,:m}$ denotes the top-left $m\times m$ submatrix. The solution to this trace maximization problem is obtained by performing an eigendecomposition on the summed covariance matrix $\sigma_{x,l}+\sigma_{y,l}$. The resulting matrix $\mathbf{U}_{l}$, whose columns are the eigenvectors sorted in descending order of their corresponding eigenvalues, is the optimal orthogonal transformation $\mathbf{U}_{l}$.
By applying this rotation, we ensure that the principal components from both the real and imaginary dimensions of the keys are aligned and concentrated within the first few heads. Consequently, we can discard the RoPE components from the remaining heads in both queries and keys while preserving the most significant positional information, thereby minimizing the performance degradation.